# RESSA: Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation

### v1

@@ Abstract
Vision-Language Models (VLMs), integrating diverse information from multiple modalities, have shown remarkable success across various tasks. However, deploying VLMs, comprising large-scale vision and language models poses challenges in resource-constrained scenarios. While pruning followed by finetuning offers a potential solution to maintain performance with smaller model sizes, its application to VLMs remains relatively unexplored, presenting two main questions: how to distribute sparsity across different modality-specific models, and how to repair the performance of pruned sparse VLMs. To answer the first question, we conducted preliminary studies on VLM pruning and found that pruning vision models and language models with the same sparsity ratios contribute to nearly optimal performance. For the second question, unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair. Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge of weights merging arises due to the incompatibility of dense LoRA modules with sparse models that destroy the sparsity of pruned models. To tackle these challenges, we propose to Repair Sparse Vision-Language Models via Sparse Cross-modality Adaptation (RESSA). RESSA utilizes cross-modality finetuning to enhance task-specific performance and facilitate knowledge distillation from original dense models. Additionally, we introduce SparseLoRA, which applies sparsity directly to LoRA weights, enabling seamless integration with sparse models. Our experimental results validate the effectiveness of RESSA, showcasing significant enhancements, such as an 11.3% improvement under 2:4 sparsity and a remarkable 47.6% enhancement under unstructured 70% sparsity. Code will be released at: https://github.com/Shwai-He/RESSA.
@@ 1 Introduction
While pruning followed by finetuning has significantly improved the efficiency of vision models Frankle & Carbin (2019); Kusupati et al. (2020); Lee et al. (2019) and language models Chen et al. (2020); Sun et al. (2023a); Frantar & Alistarh (2023), the realm of Vision-Language Models (VLMs) remains relatively unexplored in terms of model pruning, prompting following questions: how to distribute sparsity ratios between different modality-specific models and how to repair the performance of prune sparse VLMs.
For the first question, we conducted empirical studies focusing on pruning modality-specific models. Specifically, we maintained a constant summation of sparsity ratios for language models and vision models and then experimented with various combinations of sparsity ratios. Surprisingly, we discovered that pruning the vision model and language model with the same sparsity ratios contributes to nearly optimal performance. However, even so, pruning both vision models and language models simultaneously would significantly degrade performance, especially for structured sparsity patterns (e.g., N:M sparsity Zhang et al. (2022); andYukun Ma et al. (2021)). While this degradation can be mitigated by only pruning language models, there still exists a significant performance gap compared to the full model, demonstrating the importance of post-pruning repairing.
Unlike finetuning unimodal sparse models, sparse VLMs involve cross-modality interactions, requiring specialized techniques for post-pruning performance repair. Moreover, while parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, a significant challenge arises due to the incompatibility of dense LoRA modules with sparse models: merging LoRA modules with sparse models would destroy the sparse pattern while maintaining LoRA modules would cause extra latency and slow down the inference speed.
In response to the above challenges, we propose Repair Sparse Vision-Language Models via Sparse Cross-Modality Adaptation (RESSA). Specifically, we define two repairing objectives: enhancing task-specific performance and minimizing divergence from the original dense model. The former objective guides the model to improve performance, while the latter facilitates knowledge transfer from dense models to pruned models. To address the incompatibility issue of LoRA, we introduce SparseLoRA finetuning, which utilizes binary masks on LoRA weights, allowing seamless integration with pruned weights. SparseLoRA is incorporated into both the vision and language model to ensure cross-modality finetuning.
Extensive experiments showcase the effectiveness of our proposed RESSA in repairing the performance of pruned sparse VLMs. For instance, as illustrated in Figure 1, RESSA boosts the performance by 13.1% for InstructBLIP-Vicuana-7B with 2:4 sparsity. In summary, our contributions are threefold:
@@
Following the empirical studies, we propose RESSA that repairs the pruned models via cross-modality adaptation and SparseLoRA finetuning.
Extensive experiments demonstrate the effectiveness and universality of RESSA across various VLMs and tasks.
@@ 2 Related Work
Model Pruning for Large Language Models. While large vision and language models have shown promising advancements, their massive parameter sizes present challenges for practical deployment Ma et al. (2023); Wang et al. (2020). To mitigate this, model pruning techniques have been introduced to remove redundant weights or structures Han et al. (2016); Alvarez & Salzmann (2016). The primary aim of model pruning is to minimize the disparity between models before and after pruning Liu et al. (2021); He & Xiao (2024); Frantar & Alistarh (2023). Various metrics, such as magnitude, gradient Yi-Lin Sung (2024), and activation Sun et al. (2023a), have been proposed to identify unimportant weights. However, pruning without finetuning often leads to a performance drop. Zhang et al. (2024) utilize reconstruction errors-based metrics to update the weights. Other than the disparity between sparse models and dense models, our method also considers the task-specific objective of repairing sparse models and knowledge distillation from the original full models.
@@ 3 Preliminary Study
Similarly, in Figure 3 when employing structured N:M sparsity andYukun Ma et al. (2021); Zhang et al. (2022) (i.e., in each contiguous block of values, values must be zero), all models encounter significant performance degradation and even collapse (2:4 for pruning both vision models and language models). This situation prompts us to contemplate the question of how to repair the performance degradation caused by pruning for VLMs.
@@ 4 Methodology
In this section, we will develop a pruning pipeline, RESSA, that repairs the pruned models via cross modality adaptation and SparseLoRA finetuning, with the illustration in Figure 4.
@@ 4.1 Pruning with Few Samples
Model pruning identifies less important weights using predefined metrics Han et al. (2016), typically measuring the reconstruction errors Sun et al. (2023a); Frantar & Alistarh (2023); Zhang et al. (2024) between models before and after pruning, such as magnitude, gradient, and activation. Calculating gradients or activation requires a small calibration dataset with few samples. With predefined metric and the calibration dataset, the weights of a model is scored as follows:
@@ 4.3 Cross-modality Adaptation
where represents the KL-divergence distance and logits denotes the output logits of the model with weights . Based on the original model with weights , both and can be obtained by forwarding with and separately. This avoids hosting additional weights during training. The overall optimization objective of RESSA is:
where is a scalar weight. The overall procedure of RESSA is shown in Figure 4.
@@ 5 Experimental Setup
Evaluation Datasets and Metrics. We evaluate the zero-shot ability of BLIP-2 and InstructBLIP on various datasets after pruning. We use VQAv2 Goyal et al. (2016), OK-VQA Marino et al. (2019), and GQA Hudson & Manning (2019) for visual question answering, NoCaps Agrawal et al. (2019) for image captioning, and Flickr30k Plummer et al. (2015) for image-text retrieval. We use CIDEr and SPICE to evaluate image captioning tasks and use TR@1 (top-1 text recall) and IR@1 (top-1 image recall) for image retrieval tasks.
Calibration and Training Datasets. Following Yi-Lin Sung (2024); Liu et al. (2023), our approach leverages a small subset of CC3M Sharma et al. (2018) for calibration and training data. The number of training samples ranges from 1k to 10k, while the number of calibration samples is 128, which has been shown to be sufficient for pruning Sun et al. (2023a); Frantar & Alistarh (2023); Zhang et al. (2024); Yi-Lin Sung (2024).
Finetuning Details. We use Adam Kingma & Ba (2015) as the optimizer with , = 0.9, 0.999. For regularization, we set the as 0.1 and grid-search the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, where we warm up the learning rate in the first 10% steps (of the total training steps). For different model scales, we select a batch size from {16, 32, 64}, and finetune 1 epoch, which is enough for convergence. We perform a grid search for the rank of SparseLoRA, considering values from {4, 8, 16, 32}. By trial and error, we found that a rank of 4 suffices for the QFormer and the vision model, while a rank of 8 optimally suits the language model.
@@ 6 Results
| Method | Sparsity | Visual Question Answering | Image Captioning | Image retrieval | Macro Avg. | ||||
| VQAv2 | OK-VQA | GQA | NoCaps | Flickr30k | |||||
| Accuracy | CIDEr | SPICE | TR@1 | IR@1 | |||||
| Full Model | 0% | 63.1 | 41.1 | 44.1 | 105.4 | 13.8 | 96.1 | 87.5 | 64.4 |
| Magnitude | 50% | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.0 |
| Gradient | 55.1 | 35.7 | 39.8 | 92.3 | 11.6 | 91.4 | 81.6 | 58.2 | |
| SparseGPT | 56.1 | 35.5 | 40.6 | 98.7 | 13.3 | 95.8 | 86.2 | 60.9 | |
| Wanda | 57.7 | 35.4 | 41.9 | 100.1 | 13.4 | 95.2 | 84.5 | 61.2 | |
| ECoFLaP | 57.5 | 36.2 | 42.1 | 99.0 | 12.5 | 95.7 | 85.8 | 61.3 | |
[…]
| RESSA | 61.2 | 39.5 | 43.5 | 106.6 | 14.1 | 96.0 | 87.2 | 64.0 | |
6.1 Main Experimental Results
[…]
| Method | VQA | NoCaps | ||
| VQAv2 | GQA | CIDEr | SPICE | |
| Full Model | 76.7 | 49.1 | 123.9 | 15.9 |
| 2:4 Sparsity | ||||
[…]
| RESSA | 68.3 | 45.4 | 119.3 | 15.5 |
| 4:8 Sparsity | ||||
[…]
| RESSA | 71.4 | 46.5 | 121.6 | 15.6 |
N:M Sparsity. In addition to unstructured sparsity, we also conduct experiments on N: M sparsity andYukun Ma et al. (2021); Zhang et al. (2022), which can be applied to specific GPU cores and has more practical applications Mishra et al. (2021). Compared to unstructured pruning, structured pruning causes a more significant performance drop and requires more extensive repairs. Under more structured patterns, RESSA recovers more performance, achieving a 10.5% improvement for 2:4 sparsity compared to 3.4% for unstructured sparsity. After repairing, all structured pruned models maintain over 90% of the performance of the original models, demonstrating the universality and effectiveness of RESSA.
Pruning Single Modality. Language models typically have much larger parameter sizes compared to the vision models in vision-language models, Li et al. (2023); Dai et al. (2023) (e.g., 7B for Vicuna Chiang et al. (2023) vs. 1.3B for EVA-ViT in parameters Sun et al. (2023b)). As a result, the efficiency bottleneck primarily stems from the language model component. This prompted us to investigate the impact of solely pruning language models in VLMs, with experimental results presented in Table 2. With additional parameters in the vision model component, RESSA repairs InstructBLIP with significant improvement (e.g., from 69.0 to 71.6 on VQAv2), achieving performance comparable to the Full Model. Therefore, pruning language models only is an effective way to maintain performance and efficiency.
@@ 6.1 Main Experimental Results
| InstructBLIP Vicuna-7B | RESSA | |||||||
@@
| RESSA | ||||||||
6.2 Ablation Studies
| Method | Sparsity | VQAv2 | OK-VQA | GQA |
| Wanda | 4:8 | 54.4 | 39.5 | 42.4 |
@@ 6.2 Ablation Studies
| RESSA-L | 66.4 | 46.7 | 44.3 | |
| RESSA-Q | 62.5 | 40.2 | 43.3 | |
@@
| RESSA-VLQ | 69.5 | 47.4 | 45.8 |
Cross-Modality Adaptation. To assess the efficacy of cross-modality adaptation, we integrate SparseLoRA into various models. Specifically, we denote the QFormer, vision model, and language model as ”Q”, ”V” and ”L” respectively. Different insertion settings are represented using combinations of these notations (e.g., “QLV” and “LV”). As shown in Table 4, finetuning within a single model contributes to repairing the performance, and finetuning models across different modalities further enhances the performance. Moreover, adjusting the QFormer, which aligns the dense vision model and language model, improves the cross-modality ability of sparse VLMs. In summary, jointly finetuning modality-specific models and the learnable interface contribute most to performance repair.
@@
SparseLoRA Finetuning. LoRA weights cannot be merged with pruned weights, as this would disrupt the sparse pattern. Consequently, the presence of remaining LoRA modules leads to latency and slows down inference significantly Dery et al. (2024); Rücklé et al. (2021). To address this issue, SparseLoRA aims to resolve the unmerged weights of LoRA and eliminate the latency caused by LoRA modules. Table 5 compares the performance of LoRA with SparseLoRA for VLMs with sparse language models. Remarkably, SparseLoRA finetuning achieves improved performance with fewer trainable parameters, consistent with findings from He et al. (2022).
@@
Finetuning Objectives. We further investigate the impact of the proposed finetuning objectives on BLIP-2-. In Table 6, we consider three finetuning objectives: , , and . guides the task-specific performance while guides knowledge transferring from the original full model to the pruned dense model. either minimizing or improves the performance. In addition, jointly minimizing and helps the pruned models further recover performance.
Sparsity. To assess the effectiveness of RESSA across a broader range of sparsity ratios, we experimented on InstructBLIP-Vicuna-7B with unstructured sparsity ratios ranging from 40% to 80%. As the sparse ratio exceeded 50%, the performance of pruned models began to deteriorate, eventually collapsing when , highlighting the necessity of repairing. In such scenarios, RESSA significantly improved performance, particularly for higher sparsity ratios, achieving a recovery of 47.61% of scores at and 32.65% at .
Training samples. RESSA utilizes calibration datasets for pruning and retraining. We conducted experiments to explore the impact of the number of training samples. Specifically, we randomly sampled ( = 0, 100, 1k, 10k, 100k) training data points from CC3M Sharma et al. (2018) to finetune with 50% sparsity and report the average performance of visual question answering and image caption. As shown in Figure 7, we found that finetuning pruned VLMs with few-shot samples (i.e., 100) can improve performance by a substantial margin. Further finetuning with 10k training data points resulted in a significant boost in cross-modality ability. This suggests that a small amount of data is sufficient to repair the pruned vision-language models, leveraging the knowledge and capabilities acquired during pretraining Zhou et al. (2023). When 10k, the model’s capability continues to improve with more training data and gradually becomes saturated.

### v2

# Rethinking Pruning for Vision-Language Models: Strategies for Effective Sparsity and Performance Restoration

@@ Abstract
Vision-Language Models (VLMs) integrate information from multiple modalities and have shown remarkable success across various tasks. However, deploying large-scale VLMs in resource-constrained scenarios is challenging. Pruning followed by finetuning offers a potential solution but remains underexplored for VLMs. This study addresses two key questions: how to distribute sparsity across different modality-specific models, and how to restore the performance of pruned sparse VLMs. Our preliminary studies identified two effective pruning settings: applying the same sparsity to both vision and language models, and pruning only the language models. While LoRA finetuning aims to restore sparse models, it faces challenges due to incompatibility with sparse models, disrupting the pruned sparsity. To overcome these issues, we propose SparseLoRA, which applies sparsity directly to LoRA weights. Our experimental results demonstrate significant improvements, including an 11.3% boost under 2:4 sparsity and a 47.6% enhancement under unstructured 70% sparsity. Code is released at: https://github.com/Shwai-He/VLM-Compression.
Rethinking Pruning for Vision-Language Models: Strategies for Effective Sparsity and Performance Restoration
Shwai He1 Ang Li1 Tianlong Chen2††thanks: Corresponding author 1University of Maryland, College Park 2University of North Carolina at Chapel Hill shwaihe@umd.edu, angliece@umd.edu, tianlong@cs.unc.edu
@@ 1 Introduction
While pruning followed by finetuning has significantly improved the efficiency of vision models Frankle and Carbin (2019); Kusupati et al. (2020); Lee et al. (2019) and language models Chen et al. (2020); Sun et al. (2023a); Frantar and Alistarh (2023), the realm of Vision-Language Models (VLMs) remains relatively unexplored in terms of model pruning, prompting following questions: how to distribute sparsity ratios between different modality-specific models and how to restore the performance of prune sparse VLMs.
For the first question, we conducted empirical studies on pruning modality-specific models, experimenting with various combinations of sparsity ratios. Surprisingly, we found that applying the same sparsity ratios to both vision and language models yields nearly optimal performance. On the other hand, since language models are usually much larger than vision models, pruning only the language models offers a beneficial trade-off between performance and efficiency. However, as sparsity ratios increase, pruning significantly degrades performance, especially with structured sparsity patterns (e.g., N: M sparsity Zhang et al. (2022); andYukun Ma et al. (2021)), underscoring the importance of post-pruning restoration.
While parameter-efficient LoRA finetuning has been proposed to repair the performance of sparse models, it faces a significant challenge due to the incompatibility of dense LoRA modules with sparse models. Merging LoRA modules with sparse models would destroy the sparse pattern, while maintaining LoRA modules would introduce extra latency and slow down the inference speed. To address the incompatibility issue of LoRA, we introduce SparseLoRA finetuning, which utilizes binary masks on LoRA weights, allowing seamless integration with pruned weights.
Extensive experiments showcase the effectiveness of our proposed methods in repairing the performance of pruned sparse VLMs. For instance, as illustrated in Figure 1, SparseLoRA boosts the performance by 13.1% for InstructBLIP-Vicuana-7B with 2:4 sparsity. In summary, our contributions are threefold:
@@
We propose a pipeline involving pruning and post-finetuning with SparseLoRA to restore pruned models.
Extensive experiments validate the effectiveness and universality of SparseLoRA across various VLMs and tasks.
@@ 2 Related Work
Model Pruning for Large Language Models. While large vision and language models have shown promising advancements, their massive parameter sizes present challenges for practical deployment Ma et al. (2023); Wang et al. (2020). To mitigate this, model pruning techniques have been introduced to remove redundant weights or structures Han et al. (2016); Alvarez and Salzmann (2016). The primary aim of model pruning is to minimize the disparity between models before and after pruning Liu et al. (2021); He and Xiao (2024); Frantar and Alistarh (2023). Various metrics, such as magnitude, gradient Yi-Lin Sung (2024), and activation Sun et al. (2023a), have been proposed to identify unimportant weights. However, pruning without finetuning often leads to a performance drop. Zhang et al. (2024) utilize reconstruction errors-based metrics to update the weights. Other than the disparity between sparse models and dense models, our method also considers the task-specific objective of repairing sparse models and knowledge distillation from the original full models.
@@ 3 Preliminary Study
Similarly, in Figure 3 when employing structured N:M sparsity andYukun Ma et al. (2021); Zhang et al. (2022) (i.e., in each contiguous block of values, values must be zero), all models encounter significant performance degradation and even collapse (2:4 for pruning both vision models and language models). This situation prompts us to reflect on how to restore the pruning-caused performance degradation for VLMs.
@@ 4 Methodology
In this section, we will develop a pipeline that involves pruning and post restoration, with the illustration in Figure 4.
@@ 4.1 Pruning with Few Samples
Model pruning identifies less important weights using predefined metrics Han et al. (2016), typically measuring the reconstruction errors Sun et al. (2023a); Frantar and Alistarh (2023); Zhang et al. (2024) between models before and after pruning, such as magnitude, gradient, and activation. Calculating gradients or activation requires a small calibration dataset with few samples. With predefined metric and the calibration dataset, the weights of a model is scored as follows:
@@ 4.3 Finetuning Objectives
where represents the KL-divergence distance and logits denotes the output logits of the model with weights . Based on the original model with weights , both and can be obtained by forwarding with and separately. This avoids hosting additional weights during training. The overall optimization objective of SparseLoRA is:
where is a scalar weight. The procedure of VLM pruning and SparseLoRA is shown in Figure 4.
| Method | Sparsity | Param. | Visual Question Answering | Image Captioning | Image retrieval | Macro Avg. | ||||
| VQAv2 | OK-VQA | GQA | NoCaps | Flickr30k | ||||||
| Accuracy | CIDEr | SPICE | TR@1 | IR@1 | ||||||
| Full Model | 0% | 3.9B | 63.1 | 41.1 | 44.1 | 105.4 | 13.8 | 96.1 | 87.5 | 64.4 |
| Magnitude | 50% | 2.1B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.2 | 0.1 | 0.0 |
| Gradient | 55.1 | 35.7 | 39.8 | 92.3 | 11.6 | 91.4 | 81.6 | 58.2 | ||
| SparseGPT | 56.1 | 35.5 | 40.6 | 98.7 | 13.3 | 95.8 | 86.2 | 60.9 | ||
| Wanda | 57.7 | 35.4 | 41.9 | 100.1 | 13.4 | 95.2 | 84.5 | 61.2 | ||
| ECoFLaP | 57.5 | 36.2 | 42.1 | 99.0 | 12.5 | 95.7 | 85.8 | 61.3 | ||
[…]
| Wanda + SparseLoRA | B | 61.2 | 39.5 | 43.5 | 106.6 | 14.1 | 96.0 | 87.2 | 64.0 | |
@@ 5 Experimental Setup
Evaluation Datasets and Metrics. We evaluate the zero-shot ability of BLIP-2 and InstructBLIP on various datasets after pruning. We use VQAv2 Goyal et al. (2016), OK-VQA Marino et al. (2019), and GQA Hudson and Manning (2019) for visual question answering, NoCaps Agrawal et al. (2019) for image captioning, and Flickr30k Plummer et al. (2015) for image-text retrieval. We use CIDEr and SPICE to evaluate image captioning tasks and use TR@1 (top-1 text recall) and IR@1 (top-1 image recall) for image retrieval tasks.
Calibration and Training Datasets. Following Yi-Lin Sung (2024); Liu et al. (2023), our approach leverages a small subset of CC3M Sharma et al. (2018) for calibration and training data. The number of training samples ranges from 1k to 10k, while the number of calibration samples is 128, which has been shown to be sufficient for pruning Sun et al. (2023a); Frantar and Alistarh (2023); Zhang et al. (2024); Yi-Lin Sung (2024).
Finetuning Details We use Adam Kingma and Ba (2015) as the optimizer with , = 0.9, 0.999. For regularization, we set the as 0.1 and grid-search the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, where we warm up the learning rate in the first 10% steps (of the total training steps). For different model scales, we select a batch size from {16, 32, 64}, and finetune 1 epoch, which is enough for convergence. We perform a grid search for the rank of SparseLoRA, considering values from {4, 8, 16, 32}. By trial and error, we found that a rank of 4 suffices for the QFormer and the vision model, while a rank of 8 optimally suits the language model.
@@ 6 Results
| InstructBLIP Vicuna-7B | w/SparseLoRA | |||||||
@@
| w/SparseLoRA | ||||||||
6.1 Main Experimental Results
@@
| Method | Param. | VQA | NoCaps | ||
| VQAv2 | GQA | CIDEr | SPICE | ||
| Full Model | 7.9B | 76.7 | 49.1 | 123.9 | 15.9 |
| 2:4 Sparsity | |||||
@@
[…]
| w/SparseLoRA | 68.3 | 45.4 | 119.3 | 15.5 | |
| 4:8 Sparsity | |||||
[…]
| w/SparseLoRA | 71.4 | 46.5 | 121.6 | 15.6 | |
N:M Sparsity. In addition to unstructured sparsity, we also conduct experiments on N: M sparsity andYukun Ma et al. (2021); Zhang et al. (2022), which can be applied to specific GPU cores and has more practical applications Mishra et al. (2021). Compared to unstructured pruning, structured pruning causes a more significant performance drop and requires more extensive restores. Under more structured patterns, SparseLoRA recovers more performance, achieving a 10.5% improvement for 2:4 sparsity compared to 3.4% for unstructured sparsity. After restoring, all structured pruned models maintain over 90% of the performance of the original models, demonstrating the universality and effectiveness of SparseLoRA.
Single Model Pruning. Language models typically have much larger parameter sizes compared to the vision models in vision-language models, Li et al. (2023); Dai et al. (2023) (e.g., 7B for Vicuna Chiang et al. (2023) vs. 1.3B for EVA-ViT in parameters Sun et al. (2023b)). As a result, the efficiency bottleneck primarily stems from the language model component. This prompted us to investigate the impact of solely pruning language models in VLMs, with experimental results presented in Table 3. With additional parameters in the vision model component, SparseLoRA restores InstructBLIP with significant improvement (e.g., from 69.0 to 71.6 on VQAv2), achieving performance comparable to the Full Model. Therefore, pruning language models only is an effective way to maintain performance and efficiency.
@@
6.2 Detailed Analysis
| Method | Modality | Vision + Language | Language | ||||||
| VQAv2 | OK-VQA | GQA | Avg. | VQAv2 | OK-VQA | GQA | Avg. | ||
| Wanda | – | 61.4 | 39.5 | 42.4 | 47.8 | 63.9 | 44.5 | 43.1 | 50.5 |
| w/SparseLoRA | 64.3 | 42.6 | 45.8 | 51.0 | 66.0 | 44.7 | 43.3 | 51.3 | |
| 66.4 | 46.7 | 44.3 | 52.5 | 70.8 | 46.5 | 49.5 | 55.6 | ||
| 62.5 | 40.2 | 43.3 | 48.7 | 64.3 | 44.5 | 44.1 | 51.0 | ||
| 69.5 | 47.4 | 45.8 | 54.2 | 70.6 | 46.3 | 49.3 | 55.4 | ||
| 69.0 | 46.9 | 45.4 | 53.8 | 70.2 | 45.9 | 48.1 | 54.7 | ||
To evaluate cross-modality adaptation, we integrate SparseLoRA into various models. Specifically, we denote the QFormer, vision model, and language model as "", "", and "" respectively. Different configurations are represented using combinations of these notations (e.g., "" and ""). As shown in Table 5, for cross-modality pruning (i.e., Vision + Language), finetuning within a single model contributes to performance restoration, while finetuning models across two modalities further enhances performance. In cases of single modality pruning, finetuning the pruned model alone is sufficient for restoration. Notably, joint finetuning with the QFormer does not yield performance gains beyond finetuning the pruned models.
@@ 6.2 Detailed Analysis
SparseLoRA Finetuning Achieves Comparable Performance with LoRA. LoRA weights cannot be merged with pruned weights, as this would disrupt the sparse pattern. Consequently, the presence of remaining LoRA modules leads to latency and slows down inference significantly Dery et al. (2024); Rücklé et al. (2021). To address this issue, SparseLoRA aims to resolve the unmerged weights of LoRA and eliminate the latency caused by LoRA modules. Table 4 compares the performance of LoRA with SparseLoRA for VLMs with sparse language models. Remarkably, SparseLoRA finetuning achieves improved performance with fewer trainable parameters, consistent with findings from He et al. (2022).
The Effectiveness of Finetuning Objectives. We further investigate the impact of the proposed finetuning objectives on BLIP-2-. In Table 6, we consider three finetuning objectives: , , and . guides the task-specific performance while guides knowledge transferring from the original full model to the pruned dense model. either minimizing or improves the performance. In addition, jointly minimizing and helps the pruned models further recover performance.
Ablation Study on Sparsity. To assess the effectiveness of SparseLoRA across a broader range of sparsity ratios, we experimented on InstructBLIP-Vicuna-7B with unstructured sparsity ratios ranging from 40% to 80%. As the sparse ratio exceeded 50%, the performance of pruned models began to deteriorate, eventually collapsing when , highlighting the necessity of restoring. In such scenarios, SparseLoRA significantly improved performance, particularly for higher sparsity ratios, achieving a recovery of 47.6% of scores at and 32.7% at .
Ablation Study on Calibration Datasets. SparseLoRA utilizes calibration datasets for retraining. We conducted experiments to explore the impact of the number of training samples. Specifically, we randomly sampled ( = 0, 100, 1k, 10k, 100k) training data points from CC3M Sharma et al. (2018) to finetune with 50% sparsity and report the average performance of visual question answering and image caption. As shown in Figure 7, we found that finetuning pruned VLMs with few-shot samples (i.e., 100) can improve performance by a substantial margin. Further finetuning with 10k training data points resulted in a significant boost in cross-modality ability. This suggests that a small amount of data is sufficient to restoration the pruned vision-language models, leveraging the knowledge and capabilities acquired during pretraining Zhou et al. (2023). When 10k, the model’s capability continues to improve with more training data and gradually becomes saturated.